# On the universality of the halo mass function beyond $\Lambda$CDM cosmology

### v1

@@ 3.2 Ingredients
    where the filter scale $R$ and the mass $m$ enclosed in the sphere are related by $m=4\pi R^{3}\overline{\rho}_{0}/3$. $\overline{\rho}_{0}$ can be simply obtained via the critical density $\rho_{\rm c,0}=2.7754\times 10^{11}h^{-1}M_{\odot}h^{3}{\rm Mpc}^{-3}$ and the matter density parameter $\Omega_{\rm m,0}$ at present. Note that in computing the numerical integral, we set the lower limit to be $k_{\rm min}=2\pi/L_{\rm box}$, since in the simulations there is no power on scales larger than the box wavemode. For the upper limit we take this to be $k_{\rm max}=\pi/l_{\rm soft}$ where $l_{\rm soft}$ is the softening length. We have checked that the integral has converged to much better than sub-percent precision for these limits.
@@ 3.3 Estimate of the binned mass function
We now describe how we go about estimating the HMF ${\rm d}n(m)/{\rm d}\log m$ and the multiplicity function $\nu f(\nu)$, as given by Eq. (8), from our simulations.
@@
Owing to the fact that to correct for this bias robustly would require a suite of even higher-resolution simulations, we decided to impose more conservative approach and apply a minimum cut-off in the number of particles that we would accept haloes to be sampled by. We did not impose a single cut-off, but tailored these according to our aims. To be more specific, we adopt the a particle cut-off of more than 700 particles for the case when we study the absolute value for the mass function and its evolution with redshift, which restricts the systematic bias to below $2\%$ according to Eq. (17). By contrast, when looking at the relative comparison between our variations in the cosmological model and the fiducial model, we are more tolerant and impose a cut-off at 150 particles (below $4\%$ according to Eq. (17). This is due to the fact that, this systematic bias should be, roughly speaking, cancelled or at the very least significantly reduced, when we inspect the ratios of $\nu f(\nu)$ measured from different models. In addition, we also choose to discard the highest mass bins, removing those with fewer than 100 haloes per bin. This limits the Poisson errors to better than $10\%$.
@@
In Appendix A we investigate the effect of bin width on our binned mass functions, and find that binning leads to an overestimate of the mass function, especially towards high redshift and for the higher masses or equivalently rare peaks. As a result, we also develop a binning correction method, the details of which can be found there. However, in this work, since we primarily make use of relative differences in mass functions between different cosmologies rather than focus on the absolute value, we find that the binning correction factors also cancel in the ratios. Finally, we note that, if one wishes to accurately measure the absolute value of the halo mass function to high precision, then binning corrections should be taken into account when less than 10 bins per decade of halo mass are used.
@@
Figure 2 shows a simple comparison between the measured $\nu f(\nu)$ and the fitting functions in Sheth et al. (2001) (dotted line), Courtin et al. (2011) (dashed line), and Watson et al. (2013) (dash-dotted line). The $\nu f(\nu)$ at all 5 redshifts, i.e, $z=0.00$, $0.51$, $0.99$, $1.50$, and $2.05$ are measured from a total of 10 runs of the fiducial cosmology, amounted to a volume of $1.25\ h^{-3}\rm{Gpc}^{3}$. It can be seen that for the regime $\nu\geq 5$ that all of the data points form a similar locus. However, for the regime $\nu<5$ the data for $z=0$ are roughly 5% higher amplitude than the results for higher redshift. However for these data they exhibit a redshift-independent universality. Thus for our fiducial simulations we see that universality can only be said to be approximately true. These findings are in agreement with earlier results from Crocce et al. (2010) and Courtin et al. (2011).
@@ 3.4 Estimate of the error on the HMF
The figure clearly demonstrates that the errors depend only weakly on our choice for $N_{\rm sub}$ over the whole mass range. In what follows we choose $N_{\rm sub}=64$ to compute the errors for later analysis, which corresponds to a size of sample box of $L_{\rm box}=125\ h^{-1}\rm{Mpc}$. In addition, the estimated error agrees well with the fractional Poisson error (dashed lines) at the high-mass end where the sample size is relatively small. As the number of haloes grows towards the low-mass bins, the error becomes dominated by sample variance, which leads to an increase of statistical error over the Poisson error (Crocce et al., 2010; Smith & Marian, 2011).
@@
Figure 4 shows the correlation matrix $\mathbf{r}_{ij}$ estimated using all 10 runs in the fiducial cosmology at $z=0$, which amounts to a total volume of $1.25\ h^{-3}\rm{Gpc}^{3}$. The estimate is constructed using $N_{\rm sub}=64$ per simulation. The figure clearly shows that the estimates of low- to intermediate group-mass bins are highly covariant, i.e., the cross-correlation coefficient $\mathbf{r}_{ij}\gtrsim 0.5$ for haloes with $M\lesssim\times 10^{14}\ h^{-1}{\rm M}_{\odot}$). The correlation matrix only becomes close to diagonal towards the highest mass halo bins with $M\sim 10^{15}\ h^{-1}{\rm M}_{\odot}$. Each of the 10 runs has a similar correlation matrix but noiser, the results of which can be found in the Appendix B.
@@
3.5 Testing the universality of $\nu f(\nu)$
@@
To do this then we first measure $\nu f(\nu)$ for our fiducial suite of simulations at each epoch that we have selected as described in the previous sections. While averaging over all the simulations would give us the best estimate of the locus for each redshift, we do not do this. Instead we take only the results for our run 1 simulation, since its random phase field matches that of our variation simulations and so features that can be attributed to large-scale cosmic variance can, to some extent, be mitigated in the relative difference. At each epoch, we therefore set up a linear interpolator function of the measured fiducial locus for $\nu f(\nu)$ in log-log scale. We then measure the $\nu f(\nu)$ in the variational runs and compare these data with the corresponding value obtained from the interpolated locus. In this way we are able to determine to what extent the cosmological variations can be said to hold to the locus from the fiducial run, at that epoch, respectively.
@@
where $\Delta\nu f(\nu)$ is the difference between the value of $\nu f(\nu)$ measured in the variational run and the interpolated locus from the fiducial run. $\sigma$ is the estimated error, $n$ is the number of degrees of freedom, i.e., the number of data points minus the number of fitted parameters. In our case, we do not include any parameters in the function of the locus. The reduced chi-square is usually used to describe the goodness-of-fit, with a value of $\chi_{n}^{2}\lesssim 1$ indicating a good fit since the deviations of the data are about the same or smaller than the size of their errors.
@@ 4 Results
In this section, for reference, we present the results for ${\rm d}n/{\rm d}\log{M}$ at $z=0$ for the variational and fiducial runs. We then present the results of $\nu f(\nu)$ measured from these runs at the same epoch. Finally, we include the results of $\nu f(\nu)$ at higher redshifts to show how the universality evolves with redshift honing in on the response with changes to each of the cosmological parameters varied.
@@ 4.1 Comparison of ${\rm d}n/{\rm d}\log{M}$ at $z=0$
Figure 5, top panels, shows the absolute mass function ${\rm d}n/{\rm d}\log{M}$ estimated for the variational and fiducial runs at $z=0$. The left panels show the impact for variations in the dark energy density parameter $\Omega_{\rm DE}$ and the equation of state parameters $w_{0}$ and $w_{a}$, along with the variations due to the physical density of CDM $\omega_{\rm c}$. The right panel shows the variations with the power spectral amplitude $A_{\rm s}$, its spectral index $n_{\rm s}$ and its running $\alpha_{\rm s}$, along with the physical density of baryons $\omega_{\rm b}$. The bottom panels show the relative difference between each of the variations and the fiducial model for a given cosmological parameter. We find that, at this redshift, the largest differences are for the variation of $\Omega_{\rm DE}$ (equivalently $\Omega_{\rm m}$), which are greater than $15\%$ and span nearly the whole mass range ($2\times 10^{11}\leq M\leq 5\times 10^{14}\ h^{-1}{\rm M}_{\odot}$). For other models, the deviations are almost within $5\%$ (grey area) for $M\lesssim 8\times 10^{13}\ h^{-1}{\rm M}_{\odot}$, but beyond this mass range growing differences can be observed.
As expected, the variations of some parameters affect ${\rm d}n/{\rm d}\log{M}$ in a very similar way but with differing strength. For example, the variations with respect to $w_{0}$, $w_{a}$, $\omega_{\rm c}$, $\omega_{\rm b}$, and $A_{\rm s}$ all lead to slightly positive/negative deviations for small mass scales, and this trend gradually reverses when considering larger mass scales. By contrast, the variations with respect to $n_{\rm s}$ and $\alpha_{\rm s}$ appear to affect the mass functions on all mass-scales at a similar level. Specifically, we find the positive/negative variation of them entirely increases/decreases the halo abundance over the whole mass range. Overall, it can be seen that the mass function, when expressed in the form of ${\rm d}n/{\rm d}\log{M}$, is quite sensitive with respect to these parameters, and it is difficult to accurately quantify the effect of the varied parameters on ${\rm d}n/{\rm d}\log{M}$. This statement is strengthened when one considers the evolution of the halo abundances with redshift.
Figure 7: The same as relative difference in Fig. 6 but for more selected redshifts between $z=0$ and $6.75$. Each of the columns is for the positive (squares) and negative (triangles) variations of the same cosmological parameter, with an increase of redshift from the top down. The corresponding redshifts are shown on the right side of each row. It can be seen that an increase of redshift does not impact the approximate universality observed at $z=0$ (shown in Fig. 6). The $\nu f(\nu)$ plot for these redshifts can be found in the Appendix C (Fig. 11).
@@ 4.2 Comparison of $\nu f(\nu)$ at $z=0$
Figure 6 presents the $z=0$ mass functions recast in the form of $\nu f(\nu)$, with estimates obtained in accordance with the discussion in §3.3. The top panels show measurements from the fiducial (black circles) along with the linear-interpolation model fit to this data (dash-dotted lines). The results for the variations in cosmology are presented as the coloured circles. The other panels show the relative difference between the variational runs and the linear-interpolation model of the fiducial data. The figure clearly shows that the relative differences between the variational results and the fiducial model are almost all within ±$\pm 5\%$ over the whole range of $\nu$ except a few points in the large-$\nu$ tail. However, here the errors are dominated by the low-counts in the bins, with Poisson errors being $\Delta\nu f(\nu)/\nu f(\nu)\lesssim 10\%$, given our cut-off of 100 haloes per bin.
@@
We also note that there appears to be features in the shape of the $\nu f(\nu)$ values estimated from all of the variational simulations and these features become more accentuated as one considers the $\nu>1$ data. We have checked that the appearance of these features can be ascribed entirely to the finite volume effects and the specific phase distribution of the initial Gaussian random field in Fourier space. On averaging over all 10 of the fiducial runs the $\nu f(\nu)$ locus that we recover is much smoother and for each realisation displays different features. However, when we compare the variation models to the fiducial model, it was important not to use this averaged $\nu f(\nu)$ spectrum since in this way the sample variance effects can be better mitigated for. In Appendix C we also show that the features are suppressed when one considers the $\nu f(\nu)$$\nu$ locus as a function of redshift (see Fig. 11).
@@ 5.1 Utility of a universal FoF halo mass function
In a number of studies focused on cluster cosmology the utility of the FoF HMFs has been foregone in favour of studying haloes identified through either the spherical overdensity algorithm (Tinker et al., 2008; McClintock et al., 2019; Costanzi et al., 2019a, b). The objections mostly focus on the notion that the haloes selected by a spherical overdensity algorithm are more physically meaningful quantities since one can link the objects found to the virialisation overdensities resulting from the spherical collapse model, whereas the FoF algorithm with fixed $b=0.2$, links to haloes of varying overdensity (Courtin et al., 2011; More et al., 2011). Further, it is speculated that the cluster observable–halo mass distribution function can be more readily calibrated. These arguments can be countered by noting that it is well known that the condensations found in simulations are not spherical (for example see Courtin et al., 2011; More et al., 2011). In fact gravitational collapse leads to a distribution of prolate and oblate ellipsoids (Bardeen et al., 1986; Sheth et al., 2001; Jing & Suto, 2002). In this sense, one should consider spherical haloes as being less physically meaningful. Treating ellipsoidal haloes as spherical will lead to underestimates of the mass. Secondly, owing to the fact that the cluster observable–halo mass relation requires some detailed knowledge of galaxy formation physics and observational selection effects, there is no one-to-one correspondence between any 3D halo identified in simulations and any cluster observable. These have to be fully calibrated. With the FoF algorithm, one can at least be sure that one is not missing mass, which can happen in SO identified haloes.
Ultimately, for accurate cluster cosmology, one simply needs to carefully model either of the conditional probability density functions $P(X|M_{\rm SO})$ or $P(X|M_{\rm FoF})$, where $X$ is a given cluster observable. However, when it comes to making accurate predictions for the cluster counts, these will be obtained by convolving the above PDFs with the respective model for the mass functions. Our findings in this work suggest that the modelling of the $mn(m)$ for FoF haloes is simplified by the ability to exploit the near universality shown, especially when one considers extensions beyond the simple $\Lambda$CDM model. All of this leads us to the recommendation that it is $P(X|M_{\rm FoF})$ one should expend resources to calibrate over other choices.

### v2

@@ 3.2 Ingredients
    where the filter scale $R$ and the mass $m$ enclosed in the sphere are related by $m=4\pi R^{3}\overline{\rho}_{0}/3$. $\overline{\rho}_{0}$ can be simply obtained via the critical density $\rho_{\rm c,0}=2.7754\times 10^{11}h^{-1}M_{\odot}h^{3}{\rm Mpc}^{-3}$ and the matter density parameter $\Omega_{\rm m,0}$ at present. Note that in computing the numerical integral, we set the lower limit to be $k_{\rm min}=2\pi/L_{\rm box}$, since in the simulations there is no power on scales larger than the box wavemode. For the upper limit we take this to be $k_{\rm max}=\pi/l_{\rm soft}$ where $l_{\rm soft}$ is the softening length. We have checked that the integral has converged to much better than sub-percent precision for these limits regarding the mass range that we consider in this work. We have also explored whether the finite Fourier lattice on which the initial Gaussian random fields were generated could cause a systematic error in the mass function calculations. We found that the idealised matter variance on the lattice agrees with the continuum results to very high precision for mass scales $M\geq M_{\rm Ny}\sim 33m_{\rm p}$ (see Appendix A for details).
@@ 3.3 Estimate of the binned mass function
We now describe how we go about estimating the HMF ${\rm d}n(m)/{\rm d}\log m$ (equivalently, $mn(m)$, as defined in Eq. (8)) and the multiplicity function $\nu f(\nu)$, as given by Eq. (8), from our simulations.
@@
Owing to the fact that to correct for this bias robustly would require a suite of even higher-resolution simulations, we decided to impose more conservative approach and apply a minimum cutoff in the number of particles that we would accept haloes to be sampled by. We did not impose a single cutoff, but tailored these according to our aims. To be more specific, we adopt a particle cutoff of more than 700 particles for the case when we study the absolute value for the mass function and its evolution with redshift, which restricts the systematic bias to below $2\%$ according to Eq. (17). By contrast, when looking at the relative comparison between our variations in the cosmological model and the fiducial model, we are more tolerant and impose a cutoff at 150 particles (below $4\%$ according to Eq. (17). This is due to the fact that, this systematic bias should be, roughly speaking, cancelled or at the very least significantly reduced, when we inspect the ratios of $\nu f(\nu)$ measured from different models. In addition, we also choose to discard the highest mass bins, removing those with fewer than 100 haloes per bin. This limits the Poisson errors to better than $10\%$.
@@
In Appendix B we investigate the effect of bin width on our binned mass functions, and find that binning leads to an overestimate of the mass function, especially towards high redshift and for the higher masses or equivalently rare peaks. As a result, we also develop a binning correction method, the details of which can be found there. However, in this work, since we primarily make use of relative differences in mass functions between different cosmologies rather than focus on the absolute value, we find that the binning correction factors also cancel in the ratios. Finally, we note that, if one wishes to accurately measure the absolute value of the halo mass function to high precision, then binning corrections should be taken into account when less than 10 bins per decade of halo mass are used.
@@
Figure 3 shows a simple comparison between the measured $\nu f(\nu)$ and the fitting functions in Sheth et al. (2001) (dotted line), Courtin et al. (2011) (dashed line), and Watson et al. (2013) (dash-dotted line). The $\nu f(\nu)$ at all 5 redshifts, i.e, $z=0.00$, $0.51$, $0.99$, $1.50$, and $2.05$ are measured from a total of 10 runs of the fiducial cosmology, amounted to a volume of $1.25\ h^{-3}\rm{Gpc}^{3}$. It can be seen that for the regime $\nu\geq 5$ all of the data points form a similar locus. However, for the regime $\nu<5$ the data for $z=0$ are roughly 5% higher amplitude than the results for higher redshift, and the data for $z>0$ exhibit an approximate redshift-independent universality. Thus for our fiducial simulations we see that universality can only be said to be approximately true. These findings are in agreement with earlier results from Crocce et al. (2010) and Courtin et al. (2011).
@@ 3.4 Estimate of the error on the HMF
The figure clearly demonstrates that the errors depend only weakly on our choice for $N_{\rm sub}$ over the whole mass range. In what follows we choose $N_{\rm sub}=64$ to compute the errors for later analysis, which corresponds to a size of sample box of $L_{\rm box}=125\ h^{-1}\rm{Mpc}$. In addition, the estimated error agrees well with the fractional Poisson error (dashed lines) at the high-mass end where the sample size is relatively small. As the number of haloes grows towards the low-mass bins, the error becomes dominated by sample variance, which leads to an increase in the statistical error over the Poisson error (Crocce et al., 2010; Smith & Marian, 2011).
@@
Figure 5 shows the correlation matrix $\mathbf{r}_{ij}$ estimated using all 10 runs in the fiducial cosmology at $z=0$, which amounts to a total volume of $1.25\ h^{-3}\rm{Gpc}^{3}$. The estimate is constructed using $N_{\rm sub}=64$ per simulation. The figure clearly shows that the estimates of low- to intermediate group-mass bins are highly covariant, i.e., the cross-correlation coefficient $\mathbf{r}_{ij}\gtrsim 0.5$ for haloes with $M\lesssim\times 10^{14}\ h^{-1}{\rm M}_{\odot}$). The correlation matrix only becomes close to diagonal towards the highest mass halo bins with $M\sim 10^{15}\ h^{-1}{\rm M}_{\odot}$. Each of the 10 runs has a similar correlation matrix but with more noise, the results of which can be found in the Appendix D.
@@
3.5 Testing the universality of $\nu f(\nu)$ across cosmological variations
@@
To do this then we first measure $\nu f(\nu)$ for our fiducial suite of simulations at each epoch that we have selected as described in the previous sections. While averaging over all the simulations would give us the best estimate of the locus for each redshift, we do not do this. Instead we take only the results for our run 1 simulation, since its random phase field matches that of our variation simulations and so features that can be attributed to large-scale cosmic variance can, to some extent, be mitigated in the relative difference. At each epoch, we therefore set up a linear interpolator function of the measured fiducial locus for $\nu f(\nu)$ by simply applying the scipy.interpolate.interp1d function (Virtanen et al., 2020) in log-log scale. We then measure the $\nu f(\nu)$ in the variational runs and compare these data with the corresponding value obtained from the interpolated locus. In this way we are able to determine to what extent the cosmological variations can be said to hold to the locus from the fiducial run, at that epoch, respectively.
@@
where $\Delta\nu f(\nu)$ is the difference between the value of $\nu f(\nu)$ measured in the variational run and the interpolated locus from the fiducial run. $\sigma$ is the statistical error that can be obtained through the error on ${\rm d}n/{\rm d}\log{M}$ using Eq. (8), $n$ is the number of degrees of freedom, i.e., the number of data points minus the number of fitted parameters. In our case, we do not include any parameters in the function of the locus. The reduced chi-square is usually used to describe the goodness-of-fit, with a value of $\chi_{n}^{2}\lesssim 1$ indicating a good fit since the deviations of the data are about the same or smaller than the size of their errors.
@@ 4 Results
In this section, for reference, we present the results for ${\rm d}n/{\rm d}\log{M}$ at $z=0$ for the variational and fiducial runs. We then present the results of $\nu f(\nu)$ measured from these runs at the same epoch. Finally, we include the results of $\nu f(\nu)$ at higher redshifts to show how the universality evolves with redshift homing in on the response with changes to each of the cosmological parameters varied.
@@ 4.1 Comparison of ${\rm d}n/{\rm d}\log{M}$ at $z=0$
Figure 6, top panels, shows the absolute mass function ${\rm d}n/{\rm d}\log{M}$ estimated for the variational and fiducial runs at $z=0$. The left panels show the impact for variations in the equation of state parameters $w_{0}$ and $w_{a}$ and the dark energy density parameter $\Omega_{\rm DE}$, along with the variations due to the physical density of CDM $\omega_{\rm c}$. The right panel shows the variations with the physical density of baryons $\omega_{\rm b}$, the power spectral amplitude $A_{\rm s}$, its spectral index $n_{\rm s}$ and its running $\alpha_{\rm s}$. The bottom panels show the relative difference between each of the variations and the fiducial model for a given cosmological parameter. We find that, at this redshift, the largest differences are for the variation of $\Omega_{\rm DE}$ (equivalently a fractional variation of $16\%$ in $\Omega_{\rm m}$), which are greater than $15\%$ and span nearly the whole considered mass range ($2\times 10^{11}\leq M\leq 5\times 10^{14}\ h^{-1}{\rm M}_{\odot}$). For other models, the deviations are almost within $5\%$ (grey area) for $M\lesssim 8\times 10^{13}\ h^{-1}{\rm M}_{\odot}$, but beyond this mass range growing differences can be observed.
As expected, the variations of some parameters affect ${\rm d}n/{\rm d}\log{M}$ in a very similar way but with differing strength. For example, the variations with respect to $w_{0}$, $w_{a}$, $\omega_{\rm c}$, $\omega_{\rm b}$, and $A_{\rm s}$ all lead to slightly positive/negative deviations for small mass scales, and this trend gradually reverses when considering larger mass scales. By contrast, the variations with respect to $n_{\rm s}$ and $\alpha_{\rm s}$ appear to affect the mass functions on all mass-scales at a similar level. Specifically, we find the positive/negative variation of them entirely increases/decreases the halo abundance over the whole plotted mass range. Overall, it can be seen that the mass function, when expressed in the form of ${\rm d}n/{\rm d}\log{M}$, is quite sensitive with respect to these parameters, and it is difficult to accurately quantify the effect of the varied parameters on ${\rm d}n/{\rm d}\log{M}$. This statement is strengthened when one considers the evolution of the halo abundances with redshift.
Figure 8: The same as relative difference in Fig. 7 but for more selected redshifts between $z=0$ and $6.75$. Each of the columns is for the positive (squares) and negative (triangles) variations of the same cosmological parameter, with an increase of redshift from the top down. The corresponding redshifts are shown on the right side of each row. It can be seen that an increase of redshift does not impact the approximate universality observed at $z=0$ (shown in Fig. 7). The $\nu f(\nu)$ plot for these redshifts can be found in the Appendix E (Fig. 14).
@@ 4.2 Comparison of $\nu f(\nu)$ at $z=0$
Figure 7 presents the $z=0$ mass functions recast in the form of $\nu f(\nu)$, with estimates obtained in accordance with the discussion in §3.3. The top panels show measurements from the fiducial (black circles) along with the linear-interpolation model fit to this data (dash-dotted lines). The results for the variations in cosmology are presented as the coloured circles. The other panels show the relative difference between the variational runs and the linear-interpolation model of the fiducial data. The figure clearly shows that the relative differences between the variational results and the fiducial model are almost all within ±$\pm 5\%$ over the whole range of $\nu$ except a few points in the large-$\nu$ tail. However, here the errors are dominated by the low-counts in the bins, with Poisson errors being $\Delta\nu f(\nu)/\nu f(\nu)\lesssim 10\%$, given our cutoff of 100 haloes per bin.
@@
We also note that there appears to be features in the shape of the $\nu f(\nu)$ values estimated from all of the variational simulations and these features become more accentuated as one considers the $\nu>1$ data. We have checked that the appearance of these features can be ascribed entirely to the finite volume effects and the specific phase distribution of the initial Gaussian random field in Fourier space. On averaging over all 10 of the fiducial runs the $\nu f(\nu)$ locus that we recover is much smoother. Additionally, each individual realisation exhibits different features (see Fig. 12), which may be one way in which cosmic variance manifests. However, when we compare the variation models to the fiducial model, it was important not to use this averaged $\nu f(\nu)$ spectrum since in this way the sample variance effects can be better mitigated for. In Appendix E we also show that the features are suppressed when one considers the $\nu f(\nu)$$\nu$ locus as a function of redshift (see Fig. 14).
@@ 5.1 Utility of a universal FoF halo mass function
In a number of studies focused on cluster cosmology the utility of the FoF HMFs has been foregone in favour of studying haloes identified through either the spherical overdensity algorithm (Tinker et al., 2008; McClintock et al., 2019; Costanzi et al., 2019a, b; Bocquet et al., 2020). The objections mostly focus on the notion that the haloes selected by a spherical overdensity algorithm are more physically meaningful quantities since one can link the objects found to the virialisation overdensities resulting from the spherical collapse model, whereas the FoF algorithm with fixed $b=0.2$, links to haloes of varying overdensity (Courtin et al., 2011; More et al., 2011). Further, it is speculated that the cluster observable–halo mass distribution function can be more readily calibrated. These arguments can be countered by noting that it is well known that the condensations found in simulations are not spherical (for example see Courtin et al., 2011; More et al., 2011). In fact gravitational collapse leads to a distribution of prolate and oblate ellipsoids (Bardeen et al., 1986; Sheth et al., 2001; Jing & Suto, 2002). In this sense, one should consider spherical haloes as being less physically meaningful. Treating ellipsoidal haloes as spherical will lead to underestimates of the mass (Despali et al., 2013). Secondly, owing to the fact that the cluster observable–halo mass relation requires some detailed knowledge of galaxy formation physics and observational selection effects, there is no one-to-one correspondence between any 3D halo identified in simulations and any cluster observable. These have to be fully calibrated. With the FoF algorithm, one can at least be sure that one is not missing mass, which can happen in SO identified haloes.
Ultimately, for accurate cluster cosmology, one simply needs to carefully model either of the conditional probability density functions $P(X|M_{\rm SO})$ or $P(X|M_{\rm FoF})$, where $X$ is a given cluster observable. However, when it comes to making accurate predictions for the cluster counts, these will be obtained by convolving the above PDFs with the respective model for the mass functions. Our findings in this work suggest that the modelling of the $mn(m)$ for FoF haloes is simplified by the ability to exploit the near universality shown, especially when one considers extensions beyond the simple $\Lambda$CDM model. All of this leads us to the recommendation that it is $P(X|M_{\rm FoF})$ one could consider expending resources to calibrate. Additionally, we do not think of FoF haloes as the only path to a universal mass function, what we want to do is to provide motivation for finding a good candidate that can be used as a base model around which model variations are minimal, and so making the emulation more accurate and giving potential wider parameter coverage. We believe that FoF and SO haloes have both their advantages and disadvantages, it would be better that people can have more options of halo definitions to choose according to their research aims.
5.2 Parameter space coverage